[The following information applies to the questions displayed below.] Ramirez Company installs a computerized manufacturing machine in its factory at the beginning of the year at a cost of $85,400. The machine's useful life is estimated at 20 years, or 402,000 units of product, with a $5,000 salvage value. During its second year, the machine produces 34,200 units of product. Determine the machine’s second-year depreciation and year end book value under the straight-line method.

I cannot figure out how to get the year end book value (year 2) I've entered, $72,300, $4020, and more and they are all wrong, please help and explain.

Straight-Line Depreciation
Choose Numberator Choose Denominator = Annual Dep. Exp
Cost minus salvage estimated useful life depreciation exp
$80,400.00 20 $                      4,020.00
Year 2 Depreciation $4,020.00
Year end book value
(year 2)
Answer #1

since it is straight line depreciation method same amount will be reduced every year.

using the formula cost of machine less salvage value divided by estimated number of total units produced we get

85400-5000=80400

= 80400/402000

=0.2( this we will multiply with actual units produced every year)

so book value after producing 34200 units will be= 85400-(0.2 x 34200)

= 85400-6840

=78560 which will be the book value after 2nd year